A standard iodine solution was standardized against a 0.4123 g primary standard As2O3, requiring 40 mL of iodine solution. What is the normality of iodine solution? MW of As2O3= 197.84


1
Expert's answer
2022-03-02T03:11:19-0500

Equivalent mas1s= of I2 used = V1×N1 =40×N1

Mass of As2O3= 0.4123 g


Now,

Equivalent Mass of I2 liberated = Equivalent mass of As2O3

40×N1="\frac{0.4123×4×1000}{197.84}"

N1 = 0.208